Assume Digby Corp. is downsizing the size of their workforce by 10% (to the nearest person) next year from various strategic initiatives. Digby   is planning to conduct exit interviews to learn more about how they can improve in processes and increase productivity. The exit interviews are estimated to cost $100 per employee in additional to normal separation costs of $5000. How much will the company pay in separation costs if these exit interviews are implemented next year?
Select: 1
$3,300,210
$150,990
$1,358,910
$366,690
HUMAN RESOURCES SUMMARY
Andrews Baldwin Chester Digby
Needed Complement 603 421 732 719
Complement 603 421 732 719
1st Shift Complement 500 324 525 447
2nd Shift Complement 103 96 207 273
Overtime% 0.0% 0.0% 0.0% 0.0%
Turnover Rate 6.9% 7.9% 5.9% 10.0%
New Employees 42 33 43 72
Separated Employees 12 195 3 24
Recruiting Spend $5,000 $2,500 $5,000 $0
Training Hours 80 40 80 0
Productivity Index 107.3% 116.1% 128.5% 100.0%
Recruiting Cost $249 $117 $259 $72
Separation Cost $60 $976 $14 $122
Training Cost $965 $337 $1,171 $0
Total HR Admin Cost $1,274 $1,429 $1,445 $194
Labor Contract Next Year
Wages $31.04 $31.04 $31.04 $31.04
Benefits 2,500 2,500 2,500 2,500
Profit Sharing 2.0% 2.0% 2.0% 2.0%
Annual Raise 5.0% 5.0% 5.0% 5.0%
Starting Negotiation Position
Wages
Benefits
Profit Sharing
Annual Raise
Ceiling Negotiation Position
Wages
Benefits
Profit Sharing
Annual Raise
Adjusted Labor Demands
Wages
Benefits
Profit Sharing
Annual Raise
Answer #1

Present workforce of digby=719
downsizing amount=719*10%=71.9
New workforce=719-71.9=647.10
Seperation costs=(100+5000)*new work force
=5100*647.10=3300210
it is option A